# On Supports for Graphs of Bounded Genus [^1] Parts of this work appeared in “On Hypergraph Supports” available at https://arxiv.org/abs/2303.16515, which has now been split into two papers. This is the first paper in the series, and the second paper is available at https://arxiv.org/abs/2504.05039v2

### v2

@@ 1 Introduction
We introduce the notion of hypergraphs defined by connected subgraphs of a host graph. This yields a cleaner combinatorial model to study the geometric hypergraphs studied in [RR18]. To see this, consider an arrangement of disks in the plane. The dual arrangement graph $G$ corresponding to the disks is a plane graph that has a vertex for each cell of the arrangement, and two vertices in $G$ are adjacent if and only if the corresponding cells are separated by the boundary of a disk. Each disk then defines a connected subgraph of $G$ the subgraph induced on the cells contained in that disk. The subgraphs defined by any pair of disks are non-piercing in the graph-theoretic sense. That is, if $G$ is the dual arrangement graph of a set of disks, and $A$ and $B$ are two subgraphs of $G$ corresponding to disks $R_{A}$ and $R_{B}$ respectively, then the induced subgraphs on $V(A)\setminus V(B)$ and $V(B)\setminus V(A)$ are connected subgraphs of $G$. If we are given a set $P$ of points and a set $D$ of disks, this defines a hypergraph $(P,\mathcal{D})$ on $P$ where each $D\in\mathcal{D}$ defines a hyperedge $D\cap P$. For each cell containing a point of $P$, we choose a representative point of $P$ from this cell as its representative vertex, and assign it a color $blue$. Otherwise, the vertex corresponding to the cell is colored red. A planar support for the hypergraph defined by blue vertices and disks then easily yields a planar support for the hypergraph $(P,\mathcal{D})$. See Figure 1 for the construction of a graph and a support. This approach enables the study of geometric hypergraphs entirely within a graph-theoretic framework.
@@ 1.1 Problems studied
{problem}
We also study the question of construction of a support for the dual hypergraph defined on the same input. The dual hypergraph (,$(\mathcal{H},\{\mathcal{H}_{b}\}_{b\in\mathbf{b}(V)})$ is the hypergraph on $\mathcal{H}$, where each $b\in\mathbf{b}(V)$ defines a hyperedge $\mathcal{H}_{b}=\{H\in\mathcal{H}:b\in V(H)\}$ consisting of all $H\in\mathcal{H}$ containing $b$. A support for the dual hypergraph is called a dual support. For the dual support, observe that we can assume $\mathbf{b}(V)=V$, as such a support is also a support when $\mathbf{b}(V)$ is a proper subset of $V$.
{problem}
We now define intersection hypergraphs that are a common generalization of primal and dual hypergraphs. Let $G=(V,E)$ be a graph and let $\mathcal{H}$ and $\mathcal{K}$ be two families of connected subgraphs of $G$. An intersection hypergraph is the hypergraph (,$(\mathcal{H},\{\mathcal{H}_{K}\}_{K\in\mathcal{K}})$, i.e., a hypergraph on $\mathcal{H}$ such that each $K\in\mathcal{K}$ defines a hyperedge consisting of the subgraphs in $\mathcal{H}$ intersecting at a vertex of $K$. As in the dual setting, choosing a subset of $\mathcal{H}$ as terminals does not add any additional complication, and hence we define the intersection hypergraph on all of $\mathcal{H}$. Setting either $\mathcal{H}$ or $\mathcal{K}$ as singleton vertices yields respectively, the primal and dual hypergraphs. An intersection support is a support for the intersection hypergraph.
{problem}
@@
Figure 2: (a) and (b): Primal and Dual hypergraphs on the graph system $(G,\mathcal{H})$ where $G$ is the graph on vertices $\{a,b\ldots,f\}$, and =$\mathcal{H}=\{H_{1},\ldots H_{4}\}$ with $H_{1}=\{a,b,c,d\}$, $H_{2}=\{c,d,e\}$, $H_{3}=\{a,b,f,e\}$, $H_{4}=\{a,b,c,e\}$; (c): Intersection hypergraph for $(G,\mathcal{H}^{\prime},\mathcal{K})$ with =$\mathcal{H}^{\prime}=\{H^{\prime}_{1},\ldots,H^{\prime}_{4}\}$ and $\mathcal{K}=\{K_{1},\ldots,K_{3}\}$ where $H^{\prime}_{1}=\{a,b,c\}$, $H^{\prime}_{2}=\{a,b,c,e\}$, $H^{\prime}_{3}=\{c,e,f\}$, $H^{\prime}_{4}=\{b,e,d\}$, and $K_{1}=\{c,d\}$, $K_{2}=\{b,f\}$, $K_{3}=\{e,d\}$. Figures in (d)-(f) show the respective support graphs.
@@ 1.2 Related Work
As stated earlier, the existence of supports and related notions have been used in the analysis of algorithms under a local-search framework for packing and covering problems on geometric hypergraphs. Chan and Har-Peled [ChanH12] gave a PTAS for the Independent Set problem for an arrangement of pseudodisks in the plane. Mustafa and Ray [mustafa2010improved] gave a PTAS for the Hitting Set problem for a set of points and pseudodisks in the plane (both results apply more generally to k-admissible regions  [^5] A set of connected bounded regions $\mathcal{R}$ in the plane, each of whose boundary is a simple Jordan curve is $k$-admissible (for even $k$) if for any $R,R^{\prime}\in\mathcal{R}$, $R\setminus R^{\prime}$ and $R^{\prime}\setminus R$ are connected, their boundaries are in general position, and intersect each other at most $k$ times. If $k=2$, the regions are called pseudodisks.[whitesides1990k] ). The results of [ChanH12, mustafa2010improved] were extended by Basu-Roy et al., [BasuRoy2018] who used the same framework to design a PTAS for the Set Cover, and Dominating Set problem for the intersection graph of pseudodisks. Raman and Ray [RR18] showed the existence of a planar support for intersection hypergraphs of non-piercing regions (see Section 7 for the definition), and their result gave a unified analysis for all packing and covering problems described above. The local-search framework has also found applications in variants of the art gallery/terrain guarding problem [krohn2014guarding, DBLP:conf/wads/BandyapadhyayR17], where the geometric objects are defined implicitly as the visibility region of a guard, and the regions to be guarded are continuous regions.
@@ 2 Preliminaries
A graph system $(G,\mathcal{H})$ with $\mathcal{H}$ a collection of subgraphs of $G$ is non-piercing if each $H\in\mathcal{H}$ is connected and for any two subgraphs $H,H^{\prime}\in\mathcal{H}$, $H\setminus H^{\prime}$ induces a connected subgraph of $G$.
@@ 3.1 Existence of Supports
Let $(G,\mathcal{H})$ be a cross-free graph system of genus $g$. For any coloring $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$, there exists a primal support $Q$ of genus at most $g$ on $\mathbf{b}(V)$. If a cross-free embedding of $G$ with respect to $\mathcal{H}$ is given, then $Q$ can be constructed in $O(mn^{5})$ time where $m=|\mathcal{H}|$, and $n=|V(G)|$.
@@
While the three results above are algorithmic, we are unable at this point to prove that the algorithms to construct the dual support and the intersection support run in polynomial time, and we leave this as an open problem.
@@ 4 Non-piercing and Cross-free Systems
Each region $R\in\mathcal{R}$ induces a connected subgraph $G_{R}$ of $G$ consisting of vertices corresponding to cells in $R$. Moreover, it is easy to see that if $\mathcal{R}$ is non-piercing, and $\mathcal{R}_{G}$ is the collection of induced subgraphs of $G$ each corresponding to a region in $\mathcal{R}$, then the subgraphs in $\mathcal{R}_{G}$ are non-piercing subgraphs of $G$.
@@ 5 Vertex Bypassing
The basic tool we use in the construction of supports is the Vertex Bypassing (VB(.)(.)( . )) operation. Given a cross-free embedding of $G$ with respect to $\mathcal{H}$ and a vertex $v\in V(G)$, $\textsc{VB}(v)$ removes $v$ and modifies the neighborhood around $v$ so that the resulting system remains cross-free and is in some sense is simpler than the original system.
Definition 19 (VB$(v)$).
@@
In order to do so, we introduce the notion of $abab$-free hypergraphs. An equivalent notion, namely $ABAB$-free hypergraphs was studied by Ackerman et al.  [ackerman2020coloring], where the elements of the hypergraph are placed in a linear order instead of a cyclic order.
Definition 20 ($abab$-free).
An abstract hypergraph $(X,\mathcal{H})$ is said to be $abab$-free if the elements of $X$ can be arranged on a cycle $C$ such that for any $H,H^{\prime}\in\mathcal{H}$ there does not exist distinct $x_{1},x_{2},x_{3},x_{4}\in X$ in this cyclic order around $C$ such that $x_{1},x_{3}\in H\setminus H^{\prime}$, and $x_{2},x_{4}\in H^{\prime}\setminus H$.
The relation between cross-free graph systems and $abab$-free systems is the following: If $G$ is embedded so that it is cross-free with respect to $\mathcal{H}$, then for any vertex $v$, on applying $\textsc{VB}(v)$, the system $(C,\mathcal{H}^{\prime}_{v})$ is $abab$-free, where $C$ is the cycle on the vertices $u_{0},\ldots,u_{deg(v)-1}$ subdividing the edges $\{v,v_{i}\}$.
@@
Let $(G,\mathcal{H})$ be a cross-free graph system with a cross-free embedding with respect to $\mathcal{H}$. For any vertex $v\in V(G)$, on applying $\textsc{VB}(v)$, $(C,\mathcal{H}^{\prime}_{v})$ is $abab$-free.
@@
Since the subgraphs in $\mathcal{H}$ are induced subgraphs, $\mathcal{H}_{e}\subseteq\mathcal{H}_{v}$. Let $C=(u_{0},\ldots,u_{deg(v)-1})$ be the cycle on the subdivided vertices added on applying $\textsc{VB}(v)$. By construction $\mathcal{H}_{u_{i}}=\mathcal{H}_{e}$ where $e=\{v,v_{i}\}$, and $\mathcal{H}_{e}\subseteq\mathcal{H}_{v}$. If $(C,\mathcal{H}^{\prime}_{v})$ is not $abab$-free, let $H_{1},H_{2}$ be subgraphs in $\mathcal{H}^{\prime}_{v}$ that are not $abab$-free. This implies that $H_{1}$ and $H_{2}$ are crossing at $v$. ∎
By Proposition 21 therefore, the problem of adding a set of non-intersecting chords $D$ in Step 3 of Vertex Bypassing reduces to the following: Given an $abab$-free embedding of an $abab$-free hypergraph, can we add a set of non-intersecting chords in $C$ such that each subgraph is connected? We show in the following lemma, whose proof is in Section 5.1 that we can always add such chords.
@@
Let $C$ be a cycle embedded in the plane, and let $\mathcal{K}$ be a set of $abab$-free subgraphs of $C$. Then, we can add a set $D$ of non-intersecting chords in $C$ such that each $K\in\mathcal{K}$ induces a connected subgraph of $C\cup D$. Further, the set $D$ of non-intersecting chords to add can be computed in time $O(mn^{4})$ where $m=|\mathcal{K}|$, and $n=|C|$.
@@
Let $(G,\mathcal{H})$ be a cross-free system with a cross-free embedding of $G$ with respect to $\mathcal{H}$. Suppose we apply VB$(v)$ to vertex $v\in V(G)$. Then, each subgraph $H$ in $(G^{\prime},\mathcal{H}^{\prime})$ is connected. Further, VB$(v)$ can be done in time $O(|\mathcal{H}||V(G)|^{4})$.
@@
Let $C$ be the cycle added on the subdividing vertices around vertex $v$. Since $(G,\mathcal{H})$ is cross-free, by Proposition 21, the subgraphs $\{H\cap C:H\in\mathcal{H}^{\prime}_{v}\}$ satisfy the $abab$-free property on $C$. Therefore, by Lemma 5, there is a collection $D$ of non-intersecting chords such that each subgraph in $\mathcal{H}^{\prime}_{v}$ induces a connected subgraph of $C\cup D$. Hence, each subgraph $H\in\mathcal{H}^{\prime}$ is a connected subgraph of $G^{\prime}$ since each $H\in\mathcal{H}_{v}$ is modified only in the vertices of subdivision. Since Lemma 5 guarantees that the set $D$ of non-intersecting chords to add can be computed in time $O(|\mathcal{H}||V(G)|^{4})$, it follows that VB$(v)$ can be done in $O(|\mathcal{H}||V(G)|^{4})$ time. ∎
@@
5.1 Non-blocking Chords in $abab$-free Hypergraphs
In this section, we prove Lemma 5. Let $C=(x_{0},\ldots,x_{k-1},x_{0})$ be a cycle embedded in the plane with vertices labeled in clockwise order and let $\mathcal{K}$ be a collection of $abab$-free subgraphs on $C$. Since the subgraphs in $\mathcal{K}$ may not be connected, each $K\in\mathcal{K}$ induces a collection of runs on $C$, where a run is a consecutive sequence of vertices on $C$ that are contained in $K$.
For $i,j\in\{0,\ldots,k-1\}$, let $[x_{i},x_{j}]$ denote the vertices on the arc from $i$ to $j$ in clockwise order, with indices taken $\mathrm{mod}~{}k$. Similarly, we use $(x_{i},x_{j})$ to denote the open arc, i.e., consisting of the vertices on the arc from $i$ to $j$ except $x_{i}$ and $x_{j}$. We use $(x_{i},x_{j}]$ to denote the semi-open arc consisting of sequence of vertices from $i$ to $j$, but excluding the vertex $i$.

### v3

# On Supports for Graphs of Bounded Genus [^1] A preliminary version of these results appeared in EUROCOMB’23 [raman2023hypergraph]

@@ 1 Introduction
We introduce the notion of hypergraphs defined by connected subgraphs of a host graph. This yields a cleaner combinatorial model to study the geometric hypergraphs studied in [RR18]. To see this, consider an arrangement of disks in the plane. The dual arrangement graph $G$ corresponding to the disks is a plane graph that has a vertex for each cell of the arrangement, and two vertices in $G$ are adjacent if and only if the corresponding cells are separated by the boundary of a disk. Each disk then defines a connected subgraph of $G$ - the subgraph induced on the cells contained in that disk. The subgraphs defined by any pair of disks are non-piercing in the graph-theoretic sense. That is, if $G$ is the dual arrangement graph of a set of disks, and $A$ and $B$ are two subgraphs of $G$ corresponding to disks $R_{A}$ and $R_{B}$ respectively, then the induced subgraphs on $V(A)\setminus V(B)$ and $V(B)\setminus V(A)$ are connected subgraphs of $G$. If we are given a set $P$ of points and a set $D$ of disks, this defines a hypergraph $(P,\mathcal{D})$ on $P$ where each $D\in\mathcal{D}$ defines a hyperedge $D\cap P$. For each cell containing a point of $P$, we choose a representative point of $P$ from this cell as its representative vertex, and assign it a color blueblue. Otherwise, the vertex corresponding to the cell is colored red. A planar support for the hypergraph defined by blue vertices and disks then easily yields a planar support for the hypergraph $(P,\mathcal{D})$. See Figure 1 for the construction of a graph and a support. This approach enables the study of geometric hypergraphs entirely within a graph-theoretic framework.
@@ 1.1 Problems studied
    Primal Support
    Input: A graph $G=(V,E)$ of genus $g$, $c:V\to\{\mathbf{r},\mathbf{b}\}$, and a collection $\mathcal{H}$ of connected subgraphs of $G$.
    Question: Is there a support $Q$ on $\mathbf{b}(V)$ of genus $g$, i.e., for each $H\in\mathcal{H}$, $H\cap\mathbf{b}(V)$ induces a connected subgraph in $Q$.
We also study the question of construction of a support for the dual hypergraph defined on the same input. The dual hypergraph $(\mathcal{H},\{\mathcal{H}_{b}\}_{b\in\mathbf{b}(V)})$ is the hypergraph on $\mathcal{H}$, where each $b\in\mathbf{b}(V)$ defines a hyperedge $\mathcal{H}_{b}=\{H\in\mathcal{H}:b\in V(H)\}$ consisting of all $H\in\mathcal{H}$ containing $b$. A support for the dual hypergraph is called a dual support. For the dual support, observe that we can assume $\mathbf{b}(V)=V$, as such a support is also a support when $\mathbf{b}(V)$ is a proper subset of $V$.
    Dual Support
    Input: A graph $G=(V,E)$ of genus $g$, and a collection $\mathcal{H}$ of connected subgraphs of $G$.
    Question: Is there a dual support $Q^{*}$ of genus $g$ on $\mathcal{H}$, i.e., for each $v\in V(G)$, $\mathcal{H}_{v}=\{H\in\mathcal{H}:v\in H\}$ induces a connected subgraph of $Q^{*}$.
We now define intersection hypergraphs that are a common generalization of primal and dual hypergraphs. Let $G=(V,E)$ be a graph and let $\mathcal{H}$ and $\mathcal{K}$ be two families of connected subgraphs of $G$. An intersection hypergraph is the hypergraph $(\mathcal{H},\{\mathcal{H}_{K}\}_{K\in\mathcal{K}})$, i.e., a hypergraph on $\mathcal{H}$ such that each $K\in\mathcal{K}$ defines a hyperedge consisting of the subgraphs in $\mathcal{H}$ intersecting at a vertex of $K$. As in the dual setting, choosing a subset of $\mathcal{H}$ as terminals does not add any additional complication, and hence we define the intersection hypergraph on all of $\mathcal{H}$. Setting either $\mathcal{H}$ or $\mathcal{K}$ as singleton vertices yields respectively, the primal and dual hypergraphs. An intersection support is a support for the intersection hypergraph.
    Intersection Support
    Input: A graph $G=(V,E)$ of genus $g$, and two collections $\mathcal{H}$ and $\mathcal{K}$ of connected subgraphs of $G$.
    Question: Is there an intersection support $\tilde{Q}$ on $\mathcal{H}$ of genus $g$, i.e., for each $K\in\mathcal{K}$, the set $\mathcal{H}_{K}=\{H\in\mathcal{H}:V(H)\cap V(K)\neq\emptyset\}$ induces a connected subgraph of $\tilde{Q}$.
@@
Figure 2: (a) and (b): Primal and Dual hypergraphs on the graph system $(G,\mathcal{H})$ where $G$ is the graph on vertices $\{a,b\ldots,f\}$, and $\mathcal{H}=\{H_{1},\ldots H_{4}\}$ with $H_{1}=\{a,b,c,d\}$, $H_{2}=\{c,d,e\}$, $H_{3}=\{a,b,f,e\}$, $H_{4}=\{a,b,c,e\}$; (c): Intersection hypergraph for $(G,\mathcal{H}^{\prime},\mathcal{K})$ with $\mathcal{H}^{\prime}=\{H^{\prime}_{1},\ldots,H^{\prime}_{4}\}$ and $\mathcal{K}=\{K_{1},\ldots,K_{3}\}$ where $H^{\prime}_{1}=\{a,b,c\}$, $H^{\prime}_{2}=\{a,b,c,e\}$, $H^{\prime}_{3}=\{c,e,f\}$, $H^{\prime}_{4}=\{b,e,d\}$, and $K_{1}=\{c,d\}$, $K_{2}=\{b,f\}$, $K_{3}=\{e,d\}$. Figures in (d)-(f) show the respective support graphs.
@@ 1.2 Related Work
As stated earlier, the existence of supports and related notions have been used in the analysis of algorithms under a local-search framework for packing and covering problems on geometric hypergraphs. Chan and Har-Peled [ChanH12] gave a PTAS for the Independent Set problem for an arrangement of pseudodisks in the plane. Mustafa and Ray [mustafa2010improved] gave a PTAS for the Hitting Set problem for a set of points and pseudodisks in the plane (both results apply more generally to k-admissible regions  [^5] A set of connected bounded regions $\mathcal{R}$ in the plane, each of whose boundary is a simple Jordan curve is $k$-admissible (for even $k$) if for any $^{\prime}\in\mathcal{R}, RRR\setminus R^{\prime}$ and $R^{\prime}\setminus R$ are connected, their boundaries are in general position, and intersect each other at most $k$ times. If $k=2$, the regions are called pseudodisks.[whitesides1990k] ). The results of [ChanH12, mustafa2010improved] were extended by Basu-Roy et al., [BasuRoy2018] who used the same framework to design a PTAS for the Set Cover, and Dominating Set problem for the intersection graph of pseudodisks. Raman and Ray [RR18] showed the existence of a planar support for intersection hypergraphs of non-piercing regions (see Section 7 for the definition), and their result gave a unified analysis for all packing and covering problems described above. The local-search framework has also found applications in variants of the art gallery/terrain guarding problem [krohn2014guarding, DBLP:conf/wads/BandyapadhyayR17], where the geometric objects are defined implicitly as the visibility region of a guard, and the regions to be guarded are continuous regions.
@@ 2 Preliminaries
A graph system $(G,\mathcal{H})$ with $\mathcal{H}$ a collection of subgraphs of $G$ is non-piercing if each $H\in\mathcal{H}$ is connected and for any two subgraphs $^{\prime}\in\mathcal{H}, HHH\setminus H^{\prime}$ induces a connected subgraph of $G$.
@@ 3.1 Existence of Supports
Let $(G,\mathcal{H})$ be a cross-free graph system of genus $g$. For any coloring $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$, there exists a primal support $Q$ of genus at most $g$ on $\mathbf{b}(V)$.
@@
While the three results above are algorithmic, we are unable at this point to prove that the algorithms to construct the supports run in polynomial time, and we leave this as an open problem.
@@ 4 Non-piercing and Cross-free Systems
Each region $R\in\mathcal{R}$ induces a connected subgraph $G_{R}$ of $G$ - consisting of vertices corresponding to cells in $R$. Moreover, it is easy to see that if $\mathcal{R}$ is non-piercing, and $\mathcal{R}_{G}$ is the collection of induced subgraphs of $G$ each corresponding to a region in $\mathcal{R}$, then the subgraphs in $\mathcal{R}_{G}$ are non-piercing subgraphs of $G$.
@@ 5 Vertex Bypassing
The basic tool we use in the construction of supports is the Vertex Bypassing (VB(.)(.)) operation. Given a cross-free embedding of $G$ with respect to $\mathcal{H}$ and a vertex $v\in V(G)$, $\textsc{VB}(v)$ removes $v$ and modifies the neighborhood around $v$ so that the resulting system remains cross-free and is in some sense is simpler than the original system.
Definition 19 (VB(v)(v)).
@@
In order to do so, we introduce the notion of abababab-free hypergraphs. An equivalent notion, namely $ABAB$-free hypergraphs was studied by Ackerman et al.  [ackerman2020coloring], where the elements of the hypergraph are placed in a linear order instead of a cyclic order.
Definition 20 (abababab-free).
An abstract hypergraph $(X,\mathcal{H})$ is said to be abababab-free if the elements of $X$ can be arranged on a cycle $C$ such that for any $H,H^{\prime}\in\mathcal{H}$ there does not exist distinct $x_{1},x_{2},x_{3},x_{4}\in X$ in this cyclic order around $C$ such that $x_{1},x_{3}\in H\setminus H^{\prime}$, and $x_{2},x_{4}\in H^{\prime}\setminus H$.
The relation between cross-free graph systems and abababab-free systems is the following: If $G$ is embedded so that it is cross-free with respect to $\mathcal{H}$, then for any vertex $v$, on applying $\textsc{VB}(v)$, the system $(C,\mathcal{H}^{\prime}_{v})$ is abababab-free, where $C$ is the cycle on the vertices $u_{0},\ldots,u_{deg(v)-1}$ subdividing the edges $\{v,v_{i}\}$.
@@
Let $(G,\mathcal{H})$ be a cross-free graph system with a cross-free embedding with respect to $\mathcal{H}$. For any vertex $v\in V(G)$, on applying $\textsc{VB}(v)$, $(C,\mathcal{H}^{\prime}_{v})$ is abababab-free.
@@
Since the subgraphs in $\mathcal{H}$ are induced subgraphs, $\mathcal{H}_{e}\subseteq\mathcal{H}_{v}$. Let $C=(u_{0},\ldots,u_{deg(v)-1})$ be the cycle on the subdivided vertices added on applying $\textsc{VB}(v)$. By construction $\mathcal{H}_{u_{i}}=\mathcal{H}_{e}$ where $e=\{v,v_{i}\}$, and $\mathcal{H}_{e}\subseteq\mathcal{H}_{v}$. If $(C,\mathcal{H}^{\prime}_{v})$ is not abababab-free, let $H_{1},H_{2}$ be subgraphs in $\mathcal{H}^{\prime}_{v}$ that are not abababab-free. This implies that $H_{1}$ and $H_{2}$ are crossing at $v$. ∎
By Proposition 21 therefore, the problem of adding a set of non-intersecting chords $D$ in Step 3 of Vertex Bypassing reduces to the following: Given an abababab-free embedding of an abababab-free hypergraph, can we add a set of non-intersecting chords in $C$ such that each subgraph is connected? We show in the following lemma, whose proof is in Section 5.1 that we can always add such chords.
@@
Let $C$ be a cycle embedded in the plane, and let $\mathcal{K}$ be a set of abababab-free subgraphs of $C$. Then, we can add a set $D$ of non-intersecting chords in $C$ such that each $K\in\mathcal{K}$ induces a connected subgraph of $C\cup D$. Further, the set $D$ of non-intersecting chords to add can be computed in time $O(mn^{4})$ where $m=|\mathcal{K}|$, and $n=|C|$.
@@
Let $(G,\mathcal{H})$ be a cross-free system with a cross-free embedding of $G$ with respect to $\mathcal{H}$. Suppose we apply VB(v)(v) to vertex $v\in V(G)$. Then, each subgraph $H$ in $(G^{\prime},\mathcal{H}^{\prime})$ is connected. Further, VB(v)(v) can be done in time $O(|\mathcal{H}||V(G)|^{4})$.
@@
Let $C$ be the cycle added on the subdividing vertices around vertex $v$. Since $(G,\mathcal{H})$ is cross-free, by Proposition 21, the subgraphs $\{H\cap C:H\in\mathcal{H}^{\prime}_{v}\}$ satisfy the abababab-free property on $C$. Therefore, by Lemma 5, there is a collection $D$ of non-intersecting chords such that each subgraph in $\mathcal{H}^{\prime}_{v}$ induces a connected subgraph of $C\cup D$. Hence, each subgraph $H\in\mathcal{H}^{\prime}$ is a connected subgraph of $G^{\prime}$ since each $H\in\mathcal{H}_{v}$ is modified only in the vertices of subdivision. Since Lemma 5 guarantees that the set $D$ of non-intersecting chords to add can be computed in time $O(|\mathcal{H}||V(G)|^{4})$, it follows that VB(v)(v) can be done in $O(|\mathcal{H}||V(G)|^{4})$ time. ∎
@@
5.1 Non-blocking Chords in abababab-free Hypergraphs
In this section, we prove Lemma 5. Let $C=(x_{0},\ldots,x_{k-1},x_{0})$ be a cycle embedded in the plane with vertices labeled in clockwise order and let $\mathcal{K}$ be a collection of abababab-free subgraphs on $C$. Since the subgraphs in $\mathcal{K}$ may not be connected, each $K\in\mathcal{K}$ induces a collection of runs on $C$, where a run is a consecutive sequence of vertices on $C$ that are contained in $K$.
For $i,j\in\{0,\ldots,k-1\}$, let $x_{i},x_{j}]$ denote the vertices on the arc from $i$ to $j$ in clockwise order, with indices taken $\mathrm{mod}~k$. Similarly, we use $(x_{i},x_{j})$ to denote the open arc, i.e., consisting of the vertices on the arc from $i$ to $j$ except $x_{i}$ and $x_{j}$. We use $(x_{i},x_{j}]$ to denote the semi-open arc consisting of sequence of vertices from $i$ to $j$, but excluding the vertex $i$.